Read the first four sections of Appendix A: Measurements, Types of Errors, Absolute and Relative Error and Error Propagation. Answer the following questions: (a) The smallest division on the scale of the triple-beam balance is 0.1 g. What is the error in mass om for measurements made with this balance? (b) What is the percent error of a length L 5 cm with uncertainty oL 0.05 cm?
Answer #1

(a) This is a lab based instrument. I think for 0.1g least count, we can measure upto 0.01g .

(b) Here, relative error is given as

relative error = 0.05 / 5 = 0.01

Then percentage error is 0.01 * 100 = 1 %